### v1

@@ A.1 Unconditional Diffusion Models
$q(\bm{x}_{t}|\bm{x}_{t-1})=\mathcal{N}(\bm{x}_{t};\sqrt{1-\beta_{t}}\bm{x}_{t-1},\beta_{t}\bm{I}),$
@@
$q(\bm{x}_{t}|\bm{x}_{0})=\mathcal{N}(\bm{x}_{t};\sqrt{\bar{\alpha}_{t}}\bm{x}_{0},(1-\bar{\alpha}_{t})\bm{I}),$ (14)
@@
$p_{\bm{\theta}}(\bm{x}_{t-1}|\bm{x}_{t})=\mathcal{N}(\bm{x}_{t-1};\bm{\mu}_{\bm{\theta}}(\bm{x}_{t},t),\sigma^{2}_{\bm{\theta}}(\bm{x}_{t},t)\bm{I}),$ (16)
@@ A.2 Conditional Diffusion Models
$p_{\bm{\theta}}(\bm{x}_{t-1}|\bm{x}_{t},\bm{c})=\mathcal{N}(\bm{x}_{t-1};\bm{\mu}_{\bm{\theta}}(\bm{x}_{t},t,\bm{c}),\sigma_{t}^{2}\bm{I}),$ (21)
@@ D.1 First-Order ODE Sampler
To address this limitation, UniCardio employs Denoising Diffusion Implicit Models (DDIM) (DDIM, ), a training-free first-order ordinary differential equation (ODE) sampler. DDIM reinterprets the reverse diffusion process as a deterministic mapping, allowing high-fidelity sample generation with substantially fewer steps than stochastic Langevin-based approaches. Without loss of generality, we take conditional sampling with a single condition $\bm{c}$ as an example. At each sampling step $t$, given the noisy representation $\bm{x}_{t}$ and the estimated noise $\bm{\epsilon}_{\bm{\theta}}$ predicted by the denoising network, DDIM (DDIM, ) updates the posterior sampling from timestep $t$ to timestep $t-1$ with:
@@ Appendix E Pseudo Code of Training and Sampling
Input: Trained UniCardio model $\bm{\epsilon}_{\theta}$, selected generation target modality, observed condition inputs $\bm{c}$, noise schedule $\{\bar{\alpha}_{t}\}_{t=1}^{T}$, a set of discrete timesteps $\{t\}$ linearly spaced from $T$ to 00
@@ Appendix F Additional Results
PPG Imputation UniCardio 0.1160 ±$\pm 0.0066$ 0.1144 ±$\pm 0.0064$ 0.1180 ±$\pm 0.0065$ 0.1120 ±$\pm 0.0063$
UniCardio w/o LRS 0.1461 ±$\pm 0.0063$ 0.1856 ±$\pm 0.0070$ 0.3834 ±$\pm 0.0060$
UniCardio w/o TBC 0.1093 ±$\pm 0.0062$ 0.3791 ±$\pm 0.0056$ 0.4125 ±$\pm 0.0037$
UniCardio w/o TAM 1.9365 ±$\pm 0.0198$ 60.7303 ±$\pm 0.7026$ 52.8700 ±$\pm 0.3954$
ECG Imputation UniCardio 0.1755 ±$\pm 0.0043$ 0.1700 ±$\pm 0.0047$ 0.1882 ±$\pm 0.0040$ 0.1717 ±$\pm 0.0048$
UniCardio w/o LRS 0.2350 ±$\pm 0.0039$ 0.2205 ±$\pm 0.0039$ 0.2907 ±$\pm 0.0089$
UniCardio w/o TBC 0.1783 ±$\pm 0.0045$ 0.2916 ±$\pm 0.0034$ 0.2944 ±$\pm 0.0034$
UniCardio w/o TAM 3.0258 ±$\pm 0.0321$ 9.3516 ±$\pm 0.2086$ 34.6563 ±$\pm 0.6607$
PPG-to-ECG Translation UniCardio 0.2945 ±$\pm 0.0063$ 0.2873 ±$\pm 0.0071$ 0.2871 ±$\pm 0.0068$ 0.2759 ±$\pm 0.0067$
UniCardio w/o LRS 0.3581 ±$\pm 0.0097$ 0.3397 ±$\pm 0.0054$ 0.4916 ±$\pm 0.0089$
UniCardio w/o TBC 0.2832 ±$\pm 0.0065$ 0.3537 ±$\pm 0.0049$ 0.3642 ±$\pm 0.0042$
UniCardio w/o TAM 28.4177 ±$\pm 0.0547$ 29.0761 ±$\pm 0.0410$ 20.5246 ±$\pm 0.1508$
PPG-to-BP Translation UniCardio 11.4886 ±$\pm 0.4465$ 10.6768 ±$\pm 0.4491$ 10.7759 ±$\pm 0.4433$ 10.1675 ±$\pm 0.4153$
UniCardio w/o LRS 15.6685 ±$\pm 0.4075$ 11.9641 ±$\pm 0.4083$ 13.4415 ±$\pm 0.4052$
UniCardio w/o TBC 10.7032 ±$\pm 0.4103$ 15.0380 ±$\pm 0.4539$ 20.0815 ±$\pm 0.4638$
UniCardio w/o TAM 2601.45 ±$\pm 25.5475$ 3046.11 ±$\pm 11.7325$ 467.99 ±$\pm 8.7286$
@@
Table 3: Comparison of sampling efficiency. We evaluate the time required to generate each signal segment of 4 seconds with one-card A800 GPU. NFE, the number of function evaluations for diffusion models. RMSE, the root mean squared error between generated signals and ground-truth signals.
@@
Condition $M[--,0:L]$ $M[--,L:2L]$ $M[--,2L:3L]$ $M[--,3L:4L]$
Target $M[0:L,--]$ $M[L:2L,--]$ $M[2L:3L,--]$ $M[3L:4L,--]$

### v2

@@ A.1 Unconditional Diffusion Models
$q(\bm{x}_{t}|\bm{x}_{t-1})=\mathcal{N}(\bm{x}_{t}\semicolon\;\sqrt{1-\beta_{t}}\bm{x}_{t-1},\beta_{t}\bm{I}),$
@@
$q(\bm{x}_{t}|\bm{x}_{0})=\mathcal{N}(\bm{x}_{t}\semicolon\;\sqrt{\bar{\alpha}_{t}}\bm{x}_{0},(1-\bar{\alpha}_{t})\bm{I}),$ (16)
@@
$p_{\bm{\theta}}(\bm{x}_{t-1}|\bm{x}_{t})=\mathcal{N}(\bm{x}_{t-1}\semicolon\;\bm{\mu}_{\bm{\theta}}(\bm{x}_{t},t),\sigma^{2}_{\bm{\theta}}(\bm{x}_{t},t)\bm{I}),$ (18)
@@ A.2 Conditional Diffusion Models
$p_{\bm{\theta}}(\bm{x}_{t-1}|\bm{x}_{t},\bm{c})=\mathcal{N}(\bm{x}_{t-1}\semicolon\;\bm{\mu}_{\bm{\theta}}(\bm{x}_{t},t,\bm{c}),\sigma_{t}^{2}\bm{I}),$ (23)
@@ D.1 First-Order ODE Sampler
To address this limitation, UniCardio employs Denoising Diffusion Implicit Models (DDIM) (DDIM, ), a training-free first-order ordinary differential equation (ODE) sampler. DDIM reinterprets the reverse diffusion process as a deterministic mapping, allowing high-fidelity sample generation with substantially fewer steps than stochastic Langevin-based approaches. Without loss of generality, we take conditional sampling with a single condition $\bm{c}$ as an example. At each sampling step $t$, given the noisy representation $\bm{x}_{t}$ and the estimated noise $\bm{\epsilon}_{\bm{\theta}}$ predicted by the denoising network, DDIM (DDIM, ) updates the posterior sampling from timestep $t$ to timestep t1t-1 with:
@@ Appendix E Pseudo Code of Training and Sampling
Input: Trained UniCardio model $\bm{\epsilon}_{\theta}$, selected generation target modality, observed condition inputs $\bm{c}$, noise schedule $\{\bar{\alpha}_{t}\}_{t=1}^{T}$, a set of discrete timesteps $\{t\}$ linearly spaced from $T$ to 0
@@ Appendix F Additional Results
PPG Imputation UniCardio 0.1160 $\pm 0.0066$ 0.1144 $\pm 0.0064$ 0.1180 $\pm 0.0065$ 0.1120 $\pm 0.0063$
UniCardio w/o LRS 0.1461 $\pm 0.0063$ 0.1856 $\pm 0.0070$ 0.3834 $\pm 0.0060$
UniCardio w/o TBC 0.1093 $\pm 0.0062$ 0.3791 $\pm 0.0056$ 0.4125 $\pm 0.0037$
UniCardio w/o TAM 1.9365 $\pm 0.0198$ 60.7303 $\pm 0.7026$ 52.8700 $\pm 0.3954$
ECG Imputation UniCardio 0.1755 $\pm 0.0043$ 0.1700 $\pm 0.0047$ 0.1882 $\pm 0.0040$ 0.1717 $\pm 0.0048$
UniCardio w/o LRS 0.2350 $\pm 0.0039$ 0.2205 $\pm 0.0039$ 0.2907 $\pm 0.0089$
UniCardio w/o TBC 0.1783 $\pm 0.0045$ 0.2916 $\pm 0.0034$ 0.2944 $\pm 0.0034$
UniCardio w/o TAM 3.0258 $\pm 0.0321$ 9.3516 $\pm 0.2086$ 34.6563 $\pm 0.6607$
PPG-to-ECG Translation UniCardio 0.2945 $\pm 0.0063$ 0.2873 $\pm 0.0071$ 0.2871 $\pm 0.0068$ 0.2759 $\pm 0.0067$
UniCardio w/o LRS 0.3581 $\pm 0.0097$ 0.3397 $\pm 0.0054$ 0.4916 $\pm 0.0089$
UniCardio w/o TBC 0.2832 $\pm 0.0065$ 0.3537 $\pm 0.0049$ 0.3642 $\pm 0.0042$
UniCardio w/o TAM 28.4177 $\pm 0.0547$ 29.0761 $\pm 0.0410$ 20.5246 $\pm 0.1508$
PPG-to-BP Translation UniCardio 11.4886 $\pm 0.4465$ 10.6768 $\pm 0.4491$ 10.7759 $\pm 0.4433$ 10.1675 $\pm 0.4153$
UniCardio w/o LRS 15.6685 $\pm 0.4075$ 11.9641 $\pm 0.4083$ 13.4415 $\pm 0.4052$
UniCardio w/o TBC 10.7032 $\pm 0.4103$ 15.0380 $\pm 0.4539$ 20.0815 $\pm 0.4638$
UniCardio w/o TAM 2601.45 $\pm 25.5475$ 3046.11 $\pm 11.7325$ 467.99 $\pm 8.7286$
Table 2: Quantitative results of modality-restricted baselines. Baseline-1 denotes single-modality models trained on one-condition imputation tasks. Baseline-2 denotes dual-modality models trained on both one-condition translation tasks and two-condition imputation tasks. Baseline-3 denotes tri-modality models trained on both two-condition translation tasks and three-condition imputation tasks. All models use the same architecture, training data, and training paradigm to ensure fair comparison. The quantification results are averaged by 256 independent trials. The error bars represent the standard error of the mean.
Task Method Input Signal RMSE ($\downarrow$) MAE ($\downarrow$) KS-Test ($\downarrow$)
PPG
Imputation
Baseline-1 PPG-I $0924\pm 0.0056 0.0$$0671\pm 0.0037$ $0.1032\pm 0.0033$
UniCardio PPG-I $1146\pm 0.0063 0.0$$0797\pm 0.0043$ $0.1084\pm 0.0039$
[2pt/2pt]
ECG
Imputation
Baseline-1 ECG-I $1319\pm 0.0057 0.0$$0632\pm 0.0031$ $0.1382\pm 0.0048$
UniCardio ECG-I $1756\pm 0.0042 0.0$$0750\pm 0.0026$ $0.1486\pm 0.0055$
[2pt/2pt]
BP
Imputation
Baseline-1 BP-I $3.5269\pm 0.2433$ $2.3243\pm 0.1636$ $0.0982\pm 0.0037$
UniCardio BP-I $4.3773\pm 0.2896$ $2.8140\pm 0.1960$ $0.1123\pm 0.0041$
ECG
Imputation
Baseline-2 PPG, ECG-I $0800\pm 0.0039 0.0$$0428\pm 0.0021$ $0.1430\pm 0.0040$
UniCardio PPG, ECG-I $0657\pm 0.0039 0.0$$0344\pm 0.0021$ $0.1067\pm 0.0035$
[2pt/2pt]
PPG
Imputation
Baseline-2 PPG-I, ECG $0752\pm 0.0039 0.0$$0583\pm 0.0027$ $0.1116\pm 0.0039$
UniCardio PPG-I, ECG $0608\pm 0.0031 0.0$$0471\pm 0.0021$ $0.1010\pm 0.0034$
[2pt/2pt]
PPG-to-ECG
Translation
Baseline-2 PPG $0.2587\pm 0.0067$ $0.1778\pm 0.0071$ $0.4085\pm 0.0151$
UniCardio PPG $0.2747\pm 0.0067$ $0.1937\pm 0.0070$ $0.4407\pm 0.0154$
[2pt/2pt]
ECG-to-PPG
Translation
Baseline-2 ECG $0.1539\pm 0.0051$ $0.1202\pm 0.0040$ $0.1406\pm 0.0045$
UniCardio ECG $0.1802\pm 0.0060$ $0.1390\pm 0.0047$ $0.1491\pm 0.0046$
BP
Imputation
Baseline-2 PPG, BP-I $1.9208\pm 0.1260$ $1.4118\pm 0.0688$ $0.0939\pm 0.0033$
UniCardio PPG, BP-I $1.3995\pm 0.1049$ $1.0188\pm 0.0555$ $0.0766\pm 0.0027$
[2pt/2pt]
PPG
Imputation
Baseline-2 PPG-I, BP $0610\pm 0.0027 0.0$$0482\pm 0.0020$ $0.1111\pm 0.0042$
UniCardio PPG-I, BP $0445\pm 0.0016 0.0$$349\pm 0.0012 0.0$$0915\pm 0.0033$
[2pt/2pt]
PPG-to-BP
Translation
Baseline-2 PPG $9.6240\pm 0.4190$ $7.8722\pm 0.8473$ $0.2448\pm 0.0092$
UniCardio ECG $10.1538\pm 0.4213$ $8.3721\pm 0.3889$ $0.2668\pm 0.0097$
[2pt/2pt]
BP-to-PPG
Translation
Baseline-2 PPG $0.1511\pm 0.0054$ $0.1195\pm 0.0042$ $0.1311\pm 0.0046$
UniCardio PPG $0.1471\pm 0.0054$ $0.1164\pm 0.0043$ $0.1361\pm 0.0046$
ECG
Imputation
Baseline-2 ECG-I, BP $0543\pm 0.0033 0.0$$0333\pm 0.0184$ $0.1546\pm 0.0044$
UniCardio ECG-I, BP $0400\pm 0.0030 0.0$$0247\pm 0.0018$ $0.1014\pm 0.0032$
[2pt/2pt]
BP
Imputation
Baseline-2 ECG, BP-I $2.3571\pm 0.1607$ $1.7513\pm 0.1055$ $0.1010\pm 0.0382$
UniCardio ECG, BP-I $1.6834\pm 0.1056$ $1.2637\pm 0.0735$ $0.0831\pm 0.0030$
[2pt/2pt]
ECG-to-BP
Translation
Baseline-2 ECG $7.0076\pm 0.3556$ $5.8537\pm 0.3120$ $0.2070\pm 0.0089$
UniCardio ECG $8.0747\pm 0.4052$ $6.7337\pm 0.3518$ $0.2272\pm 0.0092$
[2pt/2pt]
BP-to-ECG
Translation
Baseline-2 ECG $0.1554\pm 0.0081$ $0.1110\pm 0.0073$ $0.3050\pm 0.0143$
UniCardio ECG $0.1717\pm 0.0080$ $0.1255\pm 0.0075$ $0.3433\pm 0.0150$
PPG
Imputation
Baseline-3 PPG-I, ECG, BP $0541\pm 0.0026 0.0$$0426\pm 0.0018$ $0.1012\pm 0.0033$
UniCardio PPG-I, ECG, BP $0444\pm 0.0017 0.0$$349\pm 0.0013 0.0$$0911\pm 0.0032$
[2pt/2pt]
BP&ECG-to-PPG
Translation
Baseline-3 ECG, BP $0.1833\pm 0.0064$ $0.1458\pm 0.0052$ $0.1562\pm 0.0046$
UniCardio ECG, BP $0.1324\pm 0.0044$ $0.1048\pm 0.0036$ $0.1299\pm 0.0046$
ECG
Imputation
Baseline-3 PPG, ECG-I, BP $0598\pm 0.0033 0.0$$0324\pm 0.0018$ $0.1174\pm 0.0037$
UniCardio PPG, ECG-I, BP $0386\pm 0.0028 0.0$$0241\pm 0.0017$ $0.1017\pm 0.0033$
[2pt/2pt]
PPG&BP-to-ECG
Translation
Baseline-3 PPG, BP $0.2887\pm 0.0094$ $0.2217\pm 0.0099$ $0.4622\pm 0.0158$
UniCardio PPG, BP $0.1664\pm 0.0076$ $0.1201\pm 0.0070$ $0.3352\pm 0.0147$
BP
Imputation
Baseline-3 PPG, ECG, BP-I $1.4198\pm 0.0709$ $1.0637\pm 0.0450$ $0.0782\pm 0.0026$
UniCardio PPG, ECG, BP-I $1.0723\pm 0.0437$ $8336\pm 0.0334 0.0$$0703\pm 0.0023$
[2pt/2pt]
PPG&ECG-to-BP
Translation
Baseline-3 PPG, ECG $11.6832\pm 0.5013$ $10.2349\pm 0.4833$ $0.3107\pm 0.0112$
UniCardio PPG, ECG $6.4716\pm 0.3442$ $5.4830\pm 0.3082$ $0.2013\pm 0.0085$
Table 3: Quantification results of generated signals. The evaluation metrics include the root mean squared error (RMSE), mean absolute error (MAE), signal-to-noise ratio (SNR), and Kolmogorov-Smirnov test (KS-Test), averaged by 256 independent trials. “-N” denotes the noised signals. “-I” denotes the intermittent signals.
Task Method Input Signal RMSE ($\downarrow$) SNR ($\uparrow$) KS-Test ($\downarrow$)
PPG Denoising (15 dB SNR) Wavelet Denoising sendur2003bivariate PPG-N 0.0825 18.0851 0.0740
UniCardio PPG-N 0.0285 27.3389 0.0385
UniCardio-M PPG-N, ECG 0.0274 27.6628 0.0380
UniCardio-M PPG-N, BP 0.0257 28.2883 0.0374
UniCardio-M PPG-N, ECG, BP 0.0256 28.3034 0.0376
ECG Denoising (15 dB SNR) Wavelet Denoising sendur2003bivariate ECG-N 0.0936 16.5280 0.1267
UniCardio ECG-N 0.0289 26.6413 0.0812
UniCardio-M PPG, ECG-N 0.0279 26.9689 0.0789
UniCardio-M ECG-N, BP 0.0266 27.4222 0.0788
UniCardio-M PPG, ECG-N, BP 0.0264 27.4700 0.0789
BP Denoising (15 dB SNR) Wavelet Denoising sendur2003bivariate BP-N 3.3672 28.8702 0.0862
UniCardio BP-N 1.0573 38.9173 0.0428
UniCardio-M PPG, BP-N 0.9333 40.0531 0.0408
UniCardio-M ECG, BP-N 0.9560 39.8357 0.0409
UniCardio-M PPG, ECG, BP-N 0.8756 40.6339 0.0393
Task Method Input Signal RMSE ($\downarrow$) MAE ($\downarrow$) KS-Test ($\downarrow$)
PPG Imputation GP Algorithm gerchberg1974super ; papoulis2003new PPG-I 0.7572 0.6304 0.3785
UniCardio PPG-I 0.1146 0.0818 0.1084
UniCardio-M PPG-I, ECG 0.0607 0.0465 0.0996
UniCardio-M PPG-I, BP 0.0446 0.0349 0.0918
UniCardio-M PPG-I, ECG, BP 0.0443 0.0347 0.0907
ECG Imputation GP Algorithm gerchberg1974super ; papoulis2003new ECG-I 0.8664 0.2706 0.6150
UniCardio ECG-I 0.1722 0.0848 0.1822
UniCardio-M PPG, ECG-I 0.0659 0.0345 0.1075
UniCardio-M ECG-I, BP 0.0399 0.0246 0.1012
UniCardio-M PPG, ECG-I, BP 0.0385 0.0241 0.1024
BP Imputation GP Algorithm gerchberg1974super ; papoulis2003new BP-I 87.5992 79.0766 0.9089
UniCardio BP-I 4.3977 2.8213 0.1148
UniCardio-M PPG, BP-I 1.3994 1.0226 0.0770
UniCardio-M ECG, BP-I 1.6877 1.2678 0.0834
UniCardio-M PPG, ECG, BP-I 1.0827 0.8402 0.0697
@@
Table 4: Comparison of sampling efficiency. We evaluate the time required to generate each signal segment of 4 seconds with one-card A800 GPU. The accompanying pre-processing pipeline adds only 0.0034 seconds per segment on a standard CPU of the same server. NFE, the number of function evaluations for diffusion models. RMSE, the root mean squared error between generated signals and ground-truth signals.
@@
Condition M[,0:L]M[--,0:L] M[,L:2L]M[--,L:2L] M[,2L:3L]M[--,2L:3L] M[,3L:4L]M[--,3L:4L]
Target M[0:L,]M[0:L,--] M[L:2L,]M[L:2L,--] M[2L:3L,]M[2L:3L,--] M[3L:4L,]M[3L:4L,--]
Figure 2: Performance of BP translation tasks. UniCardio is pretrained on the training set of Cuffless BP dataset. The generated signals are evaluated using unseen data of the Cuffless BP dataset kachuee2016cuffless (left) and MIMIC dataset moody1996database (right). The quantification results are averaged by 256 independent trials. The error bars represent the standard error of the mean.